OpenStudy (anonymous):

Find a linear approximation of the function f(x)=1(1+2x)4 at a=0. I got: 1+x this is not right what did i do wrong?

OpenStudy (phi):

what exactly is your function? \[ 1(1+2x)^4 \]? what is there a 1 out front? what does a=0 mean?

OpenStudy (anonymous):

\[1/(1+2x)^4\]

OpenStudy (anonymous):

that was the problem I didn't write it right

OpenStudy (phi):

f(x)= f(a) + f'(a)(x-a) OK, what did you get for the derivative

OpenStudy (anonymous):

i got 1+x

OpenStudy (phi):

\[ f(x)= \frac{1}{(1+2x)^4} = (1+2x)^{-4} \] this is (sort of) like \(u^{n}\) you should get \( n u^{n-1} du \)

OpenStudy (phi):

let u= 1+2x

OpenStudy (phi):

the derivative is not 1+x

OpenStudy (anonymous):

im not sure how to solve this problem

OpenStudy (phi):

can you take the derivative of \[ u^{-4} \]

OpenStudy (anonymous):

is it -8(1+2x)^-5

OpenStudy (phi):

If you do not know how to take the derivative of, for example \[ d (x^2) = 2x dx \] this might help... http://www.khanacademy.org/math/calculus/differential-calculus/v/calculus--derivatives-1

OpenStudy (phi):

yes, \[ \frac{d f(x)}{dx}= \frac{-8}{(1+2x)^5} \]

OpenStudy (phi):

now you use f(x)= f(a) + f'(a)(x-a) with a=0 what is f(0) ?

OpenStudy (anonymous):

is the answer 1+-8x?

OpenStudy (phi):

what is f(0)? f(0)= 1/(1+2*0)^4 = ?

OpenStudy (anonymous):

i mean 1-8x

OpenStudy (phi):

are you asking about the final answer?

OpenStudy (anonymous):

yes

OpenStudy (phi):

you need f(0), f'(0) and (x-0) f(0)= 1 f'(0)= -8 x-0 is x plug into f(x) = f(0) + f'(0)(x-0) you get f(x) = 1 + -8*x or f(x)= 1-8x

OpenStudy (anonymous):

sweet thanks for the help